write an equation of the line containing the given point and parallel to the given line.(6,-4); 6x-5y=4
Answer by TakeishaM(13)

First you need to put your equation in slope intercept form which is y = mx + b
6x-5y=4 (subtract 6x from both sides)
-5y=4-6x (divide both sides by -5; so divide 4 by -5 and -6 by -5)
y=(-4/5)+1.2x
We know that parallel lines have the same slope so the slope of your new line will be 1.2x.
Now solve for b by using the point you were given in the problem (6,-4)
y=mx+b
-4=1.2(6)+b
-4=7.2+b (subtract 7.2 from both sides)
-11.2 = b
The answer is y=1.2x-11.2